# Adversarial Evasion Attacks Practicality in Networks: Testing the Impact of Dynamic Learning Authors’ draft for soliciting feedback: April 4, 2024

### v3

@@ V-C Problem-Space Perturbations
V-D Adaptivity to Dynamic ML
VI Investigating of The Effect of Dynamic Learning on Adversarial Attacks Against ML-NIDS
VI-A Overview
As noted in Section V-D, we believe that defence that adopt dynamic model should reduce the effectiveness of adversarial attack. We set up and conduct experiments to test the effect of continuous training on the effectiveness of adversarial attacks against ML-NIDS. As illustrated in Figure 5, our experiments test the impact of adversarial attacks before (attacking on Day n) and after re-training (attacking on Day n+1, and Day n+2). In this section, we will explore the ML model, dataset, metrics, and continuous training criteria used to test the practicality of adversarial attacks against dynamic ML-based NIDS.
VI-B Dataset
The CSE-CIC-IDS2018 [32] dataset was used in our experiments. Although only 12 percent of the network intrusion detection research utilizes this dataset, it was found that it is currently the most up-to-date, and ML models have better accuracy when it is used in training and testing [11]. Moreover, the CSE-CIC-IDS2018 dataset has a low number of duplicate and uncertain data [33]. The dataset was created by Communications Security Establishment (CSE) and the Canadian Institute for Cybersecurity (CIC) in 2018/2019. In the CSV format files, each tuple has 80 features. During the creation of the dataset, numerous scenarios were constructed, and data were collected and edited daily. The dataset contains six different classes (labels); Benign, Bot, Brute Force, DoS, Infiltration, and SQL Injection. However, the authors of [34] found that CSE-CIC-IDS-2018 suffers from numerous issues regarding feature generation and labeling. We agree with their findings as we discovered a considerable number of outlier samples in the dataset that need cleansing before the standardization and training processes for constructing more robust ML models.
@@ VI-B Dataset
VI-C The Target ML-based NIDS
As seen in Figure 6, the target ML-based NIDS consists of multiple phases. We crafted our own ML-NIDS for our experiments. However, we didn’t create a feature extraction stage, as we used the features provided by the CSV files of the CSE-CIC-IDS2018 dataset. We will explore our ML-NIDS in the next subsections.
VI-C1 Pre-Processing
@@
Moreover, we found numerous outlier samples, which are values that deviate significantly from the rest of the dataset. These atypical data points can present challenges for statistical analyses, like mean and standard deviation calculations, which are essential for standardization. Furthermore, in ML training datasets, outliers can disrupt model performance, mislead the training process, lead to non-generalizable models, and increase variance, potentially causing bias and reduced accuracy [35]. For example, as seen in Figure 7a, a minuscule number of records (4 samples out of millions) of "Bwd Pkt Len Max" Feature are in the neighborhood of 6000, which is extremely far from the rest of the values. Therefore, we created a method to exclude these outlier samples from the dataset. This technique detects and removes samples with a probability of less than 0.0005 in the cumulative distribution function and those with a probability of more than 0.9995, but only if the number of these records is less than a threshold. The Frequency of "Bwd Pkt Len Max" Feature values Before and After Data Cleansing can be seen in Figure 7. After excluding the small number of outliers, the range of values for the feature changed from 0 to 6000 to 0 to 1500, as seen in Figure 7b. This outlier-excluding method is applied to all numerical features. Afterward, We performed standardization and one-hot key encoding on the dataset to prepare it for training machine learning models.
(a) The Frequency of "Bwd Pkt Len Max" Feature Values Before Data Cleansing
(b) The Frequency of "Bwd Pkt Len Max" Feature Values After Data Cleansing
Figure 7: The Frequency of "Bwd Pkt Len Max" Feature Values Before and After Data Cleansing
@@
Figure 8: Dataset’s Features Correlation Matrix
@@
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
@@
VI-C2 Training and Testing
VI-D The Adversarial Attacks in Our Experiments
The Adversarial Robustness Toolbox (ART) library [38] was used to generate FGSM [15], Projected Gradient Descent (PGD) [39], and Basic Iterative Method (BIM) [40] adversarial attacks for our experiments. We tested the attacks by adding adversarial perturbations to the entire test set.
VI-E Continuous Training
The CSE-CIC-IDS2018 dataset is partitioned by days, and we assumed that the target ML-NIDS model retrained daily. However, in continuous learning environments, there is a risk of catastrophic forgetting. Catastrophic forgetting in machine learning refers to the phenomenon where a neural network or learning model forgets previously learned information when trained on new data or tasks [41].
To address this issue, we merge all data frames from previous days for each new day, shuffle the union set, and select a random sample with a uniform distribution. This approach overcomes catastrophic forgetting by ensuring the model retains its knowledge and adapts to new data without forgetting previously trained data. We repeated this process for three consecutive days to test the impact of continuous training on adversarial attacks. This simple yet effective technique enabled us to retrain the model without performance drops or the need for continuous fine-tuning.
VII Dynamic ML Effect Findings
Section V addressed numerous issues that pose significant obstacles to adversarial attacks against Network Intrusion Detection Systems. However, even if attackers were to overcome these challenges, the implementation of dynamic learning would reduce the effectiveness of the attacks, as seen in this section.
VII-A Attack Scenario
In this section, we will present the results of our experiments that test the effect of the continuous training aspect of the dynamic learning models on adversarial attacks. Our testing scenario, as depicted in Figure 5, assumes that attackers have somehow acquired the knowledge needed to calculate gradients and craft their adversarial attacks, despite the challenges discussed in Section V regarding obtaining information about the target.
Due to the complexity of formulating adversarial attacks against NIDS, there is a time delay (at least one or two days in our experiments) before the attackers initiate the evasion. Meanwhile, the target dynamic ML-NIDS undergoes retraining, rendering the knowledge obtained by the attackers outdated. In our experiments, we aim to test the efficiency of adversarial attacks against ML-based NIDS using up-to-date knowledge of the model (Day n) for evasion versus outdated knowledge after one day of retraining the NIDS (Day n+1) and two days (Day n+2). In other words, we attack the continuously daily trained model using the gradients captured on Day n to attack the model on Day n (up-to-date gradients), Day n+1 (one-day-old gradients), and Day n+2 (two-day-old gradients).
@@ VII-A Attack Scenario
TABLE I: Effect of Continuous Training on FGSM, PGD and BIM Attacks against the ANN, SVM and CNN NIDSs, where Day n represents attacking before re-training, Day n+1 represents attacking one day after re-training, and Day n+2 represents attacking two days after re-training.
@@
VII-B Results
In this subsection, we present the results of continuous training on different ML-based NIDSs (ANN, SVM, CNN) and their susceptibility to three different attack techniques. Carlini et al. [42] recommended using strong iterative attacks in addition to FGSM to evaluate adversarial attacks. Therefore, we used FGSM, PGD, and BIM attacks in our experiments. The performance metrics considered were Accuracy, Precision, Recall, and F1-score.
Figures 10 and 11 and Table I illustrate the impact of FGSM, PGD, and BIM attacks on three different NIDSs: Artificial Neural Network, Support Vector Machine, and Convolutional Neural Network. These attacks were conducted on various days. Specifically, ’Day n’ indicates an attack before re-training (with current knowledge of the model), ’Day n+1’ corresponds to an attack one day after re-training (with knowledge that’s one day old), and ’Day n+2’ signifies an attack after two re-training cycles (with knowledge that’s two days old).
Pre-attack, as seen in Table I (a, b, and c), the NIDSs achieved high accuracy, precision, recall, and F1-score, indicating their effectiveness in detecting intrusions. However, after the FGSM attack, the performance dropped significantly for all three models. Among them, the SVM model experienced the highest drop, with its accuracy and F1 decreasing from 0.997 and 0.997 to 0.150 and 0, respectively. On the other hand, the ANN and CNN models had a decrease in accuracy and F1 of 15 to 30 percent. The reduction in accuracy and F1-measure was particularly prominent, indicating that the attacks successfully evaded the NIDSs’ detection. Nevertheless, continuous training improved the NIDSs’ resistance to FGSM attacks. On Day n+1 and Day n+2, we observed a significant recovery in accuracy, precision, recall, and F1-score. As seen in Figures 11a and 10a, after just one day of re-training, all the models recovered an accuracy and F1-score of more than 0.9. Moreover, after one more day of re-training (Day n+2), accuracy and F1-score increased even further, except for the SVM model, where the accuracy and F1 stayed the same at 0.998, which is very high.
Similarly, as shown in Table I (d to i), the results indicate that, like the FGSM attack, the performance of all three models significantly decreased after the PGD and BIM attacks. Among them, the SVM models experienced the most pronounced decline, while the ANN and CNN models exhibited a milder decrease in accuracy and F1. Continuous training also proved to enhance the NIDSs’ resistance to PGD and BIM attacks. On Day n+1 and Day n+2, we observed a substantial recovery in accuracy, precision, recall, and F1-score compared to Day n. As depicted in Figures 11b, 10b, 11c, and 10c, after just one day of re-training, all the models achieved an accuracy and F1-score of over 0.9. For the PGD and BIM attacks, after one additional day of re-training (Day n+2), the accuracy and F1-score of the CNN model further improved. In the case of the SVM model, the accuracy and F1 remained constant at 0.998, a very high level. However, unlike FGSM, the metrics of the ANN model experienced a slight decrease (2 percent) after one more day of re-training (Day n+2) in the cases of BIM and PGD attacks. This minor decrease can be attributed to a slight instance of catastrophic forgetting.
These results suggest that attacks crafted with slightly outdated knowledge of the model might have a lower effect on ML models. That indicates that continuously training the model might introduce an additional hurdle for attackers, as they would constantly need to obtain the updated gradients of the model, which can be a complex task, especially in the NIDS domain. Moreover, it is worth mentioning that this recovery of the model’s measures occurred after just one and two re-training sessions. Furthermore, these results were achieved without adversarial training, meaning that the re-training data didn’t contain adversarial samples [20].
Figure 12: Decision Boundary Evolution using t-SNE
On another note, although the results between the three attacks used in our experiments, FGSM, PGD, and BIM, are close, they are different attacks with distinct generation algorithms and dissimilar generation times, as seen in Table II. FGSM is a one-shot attack, whereas PGD and BIM are iterative attacks. This suggests that the impact of continuous training on adversarial attacks may depend more on the ML models than on the attack methods themselves, which is why we have included all these results in the paper. Also, these results demonstrate that the impact of continuous training remains consistent across different attacks. Furthermore, the results reveal that various ML models exhibit different responses to adversarial attacks and varying recovery rates after retraining. These similarities and dissimilarities warrant further investigation in the future.
TABLE II: Generation of Adversarial Attacks Average Time
Attack Model ANN SVM CNN
FGSM 7.2 sec 140.5 sec 42.3 sec
PGD 51.2 sec 2137.0 sec 338.9 sec
BIM 40.2 sec >30 min 305.3 sec
VII-C Interpretation of the Results
Because numerical results alone don’t tell the entire story, and to gain an intuitive understanding of the data and results, visualizations are necessary. To achieve this, we combine the visualization of Voronoi-based representations of multidimensional decision boundaries in scatterplots with the visualization of distances to the multidimensional boundary of data elements [43]. Data elements are reduced from 50 dimensions to 2 dimensions using the t-SNE (t-distributed Stochastic Neighbor Embedding) algorithm [44]. In other words, we show the evolution of the optimal multidimensional decision boundaries in 2D after every re-training session, as seen in Figure 12. Moreover, we also visualize the evolution of the data distribution after every re-training, as shown in Figure 13. As expected, the decision boundaries and training data distribution are always changing and evolving, as seen in Figures 12 and 13.
So re-training changes the model, including the decision boundary and gradients, and adversarial examples tend to be near this constantly evolving boundary rather than deep within the wrong class. Therefore, attacks formulated with outdated knowledge of the model might have a reduced effect on machine learning models even without adversarial training.
VIII Discussion and Lessons Learned
This section discusses several lessons learned and insights gathered from our research.
VIII-A Dynamic Learning
  • Dynamic ML research is scarce in the machine learning-based NIDS domain. This scarcity needs attention because it is expected that concept drifts will happen in the NIDS space similar to other research domains [7]
  • The dynamic learning effect on adversarial attacks depends on numerous factors such as the model used, re-learning rate, re-learning data, etc. Our work just shows that it could reduce the effect of adversarial attacks on attacked models.
  • Systems that adapt to changes in legitimate behavior over time, such as evolving machine learning approaches, may be vulnerable to manipulation by intruders who gradually introduce their own malicious behavior into the baseline. This is known as session creep [47].
  • In continuous learning environments, a critical challenge is the potential occurrence of catastrophic forgetting, a phenomenon where machine learning models forget previously learned information when trained on new data or tasks [41]. To ensure the successful deployment of such models in real-world conditions, effective strategies for handling catastrophic forgetting must be carefully implemented.
VIII-B Network Security Datasets
  • Obtaining Labeled datasets to train even static ML-NIDS is very tough and expensive [48]. Unfortunately, dynamic ML models need more labeled datasets for continuous training, and acquiring these labeled datasets is even harder. This topic needs more research from network security and ML communities.
  • Although CSE-CIC-IDS-2018 is widely regarded as a reliable/up-to-date NIDS dataset [11], it has been discovered that it has several issues with feature generation and labeling [34]. This re-emphasises the issue that the ML and NIDS research communities should prioritize the development of more robust datasets.
VIII-C Adversarial attacks in network security vs computer vision
  • Crafting adversarial attacks in network/cybersecurity is more complex than in domains like computer vision [31]. Unlike subtly altering pixel values in image classification, adding perturbations to network packets or PE files can disrupt their network or malicious functionality. White-box attacks require extensive knowledge of the targeted ML model. This is often impractical in many domains like computer vision but even more challenging in network security due to heightened security measures [31]. Black-box attacks assume querying the model for gradients, but this may be unfeasible in Network Intrusion Detection Systems (NIDS) contexts.
  • On the other hand, if attackers manage to overcome all obstacles and create a practical problem space adversarial attack in the network/cybersecurity domain, the perturbed network elements and features will become more diverse and heterogeneous. This complexity makes detecting and defending against problem space adversarial attacks in the network/cybersecurity domain more challenging than in other domains, such as computer vision [19].
Figure 14: Adversarial Attacks Visualization
IX Conclusion and Future Work
In this paper, we extended the research discussion on the practicality of adversarial attacks against NIDS by identifying the questionable practicality prerequisites required to carry out evasion adversarial attacks against NIDS using an attack tree and by presenting a taxonomy of practicality issues related to ML-based NIDS adversarial attacks. Our attack tree and taxonomy highlight several factors that could make numerous researched adversarial attacks impractical against real-world ML-based NIDSs. It is important to note that our work does not claim that adversarial attacks won’t harm ML-based NIDSs; rather, we find that the gap between research and real-world practicality is wide and deserves to be addressed. Apruzzese et al. [3] argue that real-world attackers typically do not rely on gradients or conventional adversarial attacks techniques; instead, they employ alternative simple yet effective approaches to evade ML models. If this is the case, then the current direction of ML security research needs a significant shift to address real-world challenges adequately. We also explored the impact of adversarial attacks on dynamic ML-based NIDS, finding that continuous re-training, even without adversarial training, may limit the effect of such attacks. In future research, We plan to expand our attack taxonomy and apply it to network security applications beyond NIDS (e.g., spam detection). We will also explore how more sophisticated dynamic and continuous learning techniques will perform against adversarial attacks.
X Acknowledgement

### v5

@@ 7.2 The Catch?
While practical, Debicha et al.’s [38] attack faces several significant catches that severely limit its real-world effectiveness. The most critical limitation is the impractical data requirement; the paper assumes attackers have access to 50% of the target’s training dataset, which contradicts realistic black-box scenarios where attackers would never obtain such extensive labeled network traffic data through simple “network sniffing”. Additionally, transferability uncertainty means there’s no guarantee that adversarial instances crafted for surrogate models will successfully evade the actual target NIDS. Domain constraint limitations restrict the perturbation space by requiring maintenance of semantic relationships between network features, while the small perturbation requirement means attacks may fail against robust or ensemble-based NIDS that can detect subtle traffic anomalies.
Despite the limitations mentioned in this section, we still believe that these attacks could potentially be lethal against NIDS, but we also think that for each of those attacks in their current form, there is still a catch (or catches). That’s why we put the word practical in quotation marks.
8 Investigating of The Effect of Dynamic Learning on Adversarial Attacks Against ML-NIDS
As noted in Section 6.3, we believe that adopting a dynamic model might change the effectiveness of adversarial attacks. Thus, we conduct experiments to test the effect of continuous training on the effectiveness of adversarial attacks against ML-NIDS.
8.1 Attack Scenario
In this section, we will present the results of our experiments that test the effect of the continuous training aspect of the dynamic learning models on adversarial attacks. Our testing scenario, as depicted in Figure 9, assumes that attackers have somehow acquired the knowledge needed to calculate gradients and craft their adversarial attacks, despite the challenges discussed in Section 6 regarding obtaining information about the target.
Due to the complexity of formulating adversarial attacks against NIDS, there is a time delay (at least one or two days in our experiments) before the attackers initiate the evasion. Meanwhile, the target dynamic ML-NIDS undergoes re-training, rendering the knowledge obtained by the attackers outdated. In our experiments, we aim to test the efficiency of adversarial attacks against ML-based NIDS using up-to-date knowledge of the model (Day n) for evasion versus outdated knowledge after one day of re-training the NIDS (Day n+1) and two days (Day n+2). In other words, we attack the continuously daily trained model using the gradients captured on Day n to attack the model on Day n (up-to-date gradients), Day n+1 (one-day-old gradients), and Day n+2 (two-day-old gradients).
8.2 The Target ML-based NIDS
As seen in Figure 10, the target ML-based NIDS consists of multiple phases. We crafted our own ML-NIDS for our experiments. We will explore our ML-NIDS in the next subsections.
@@ 8.2 The Target ML-based NIDS
8.2.1 Dataset
The CSE-CIC-IDS2018 [44] dataset was used in our experiments. It was found to be the most up-to-date, and ML models achieve better accuracy when trained and tested on it [16]. Moreover, the CSE-CIC-IDS2018 dataset has a low number of duplicate and uncertain data [45]. The dataset was created by the Communications Security Establishment (CSE) and the Canadian Institute for Cybersecurity (CIC) in 2018. In the CSV format files, each tuple has 80 features. During the dataset’s creation, numerous scenarios were constructed, and data were collected and edited daily. The dataset contains six different classes (labels): Benign, Bot, Brute Force, DoS, Infiltration, and SQL Injection. However, the authors of [46] found that CSE-CIC-IDS-2018 suffers from numerous issues regarding feature generation and labeling. We agree with their findings, as we discovered a considerable number of outlier samples in the dataset that require cleansing before the standardization and training processes for constructing more robust ML models.
8.2.2 Pre-Processing
@@
Moreover, we found numerous outlier samples, which are values that deviate significantly from the rest of the dataset. These atypical data points can present challenges for statistical analyses, like mean and standard deviation calculations, which are essential for standardization. Furthermore, in ML training datasets, outliers can disrupt model performance, mislead the training process, lead to non-generalizable models, and increase variance, potentially causing bias and reduced accuracy [47]. For example, as seen in Figure 11a, a minuscule number of records (4 samples out of millions) of “Bwd Pkt Len Max” Feature are in the neighborhood of 6000, which is extremely far from the rest of the values. Therefore, we created a method to exclude these outlier samples from the dataset. This technique detects and removes samples with a probability of less than 0.0005 in the cumulative distribution function and those with a probability of more than 0.9995, but only if the number of these records is less than a threshold. The Frequency of “Bwd Pkt Len Max” Feature values Before and After Data Cleansing can be seen in Figure 11. After excluding the small number of outliers, the range of values for the feature changed from 0 to 6000 to 0 to 1500, as seen in Figure 11b. This outlier-excluding method is applied to all numerical features. Afterward, We performed standardization and one-hot key encoding on the dataset to prepare it for training machine learning models.
(a) The Frequency of “Bwd Pkt Len Max” Feature Values Before Data Cleansing
(b) The Frequency of “Bwd Pkt Len Max” Feature Values After Data Cleansing
Figure 11: The Frequency of “Bwd Pkt Len Max” Feature Values Before and After Data Cleansing
Figure 12: Dataset’s Features Correlation Matrix
@@
8.2.3 Models’ Architectures
@@
8.2.4 Continuous Training
The CSE-CIC-IDS2018 dataset is partitioned by days, and we assumed that the target ML-NIDS model was re-trained daily (periodic re-training). However, in continuous learning environments, there is a risk of catastrophic forgetting, a phenomenon where a learning model forgets previously learned information when trained on new data or tasks [50].
To mitigate this issue, we employed a rehearsal-based approach (also known as experience replay) by merging all data frames from previous days for each new day, shuffling the union set, and selecting a random sample with a uniform distribution. This technique reduces catastrophic forgetting by allowing the model to reinforce previously learned patterns while adapting to new data. We repeated this process for three consecutive days to test the impact of continuous training on adversarial attacks.
For our three-day experimental window, this strategy provided sufficient stability (retaining old knowledge) and plasticity (learning new tasks) without requiring more complex or expensive continual learning techniques, serving as a simple yet effective baseline for our experiments.
8.3 Adversarial Attacks in Our Experiments
The Adversarial Robustness Toolbox (ART) library [51] was used to generate FGSM [20], Projected Gradient Descent (PGD) [52], and Basic Iterative Method (BIM) [53] adversarial attacks. We evaluated these attacks by adding adversarial perturbations to the entire test set.
8.4 Threat Model Assumptions Justification
As previously mentioned, we selected FGSM, PGD, and BIM attacks and periodic re-training for our experiments to evaluate the effect of dynamic learning on adversarial evasion attacks. Our experimental design intentionally tests attacks with the strongest attacker capabilities (white-box access with feature-space manipulation) against the most basic form of dynamic learning (periodic re-training). This combination provides a clear baseline evaluation framework: if the most trivial continual learning approach can effectively reduce the impact of adversarial attacks operating under maximum attacker capabilities, then it establishes strong evidence for dynamic learning’s potential effectiveness against more realistic, constrained attack scenarios. Furthermore, Carlini et al. [54] specifically recommended using strong iterative attacks (PGD and BIM) in addition to FGSM when evaluating adversarial defense mechanisms, as these attacks provide comprehensive coverage of evasion techniques.
8.5 Results
Pre-attackDay nDay n+1Day n+200.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911ANNCNNSVM
@@ 8.5 Results
Pre-attackDay nDay n+1Day n+200.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+200.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+200.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+200.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911ANNCNNSVM
@@
Pre-attackDay nDay n+1Day n+200.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911ANNCNNSVM
@@
Figures 14 and 15 demonstrate that continuous re-training might reduce the effect of FGSM, PGD, and BIM attacks across three NIDS architectures (ANN, CNN, SVM). When attacks are launched on Day n using current model knowledge, all classifiers experience significant performance degradation, with SVM suffering the most severe drop. However, at Day n+1, which corresponds to an attack one day after re-training (with knowledge that’s one day old), all models recover to over 90% accuracy and F1-score. By Day n+2, which corresponds to an attack two days after re-training (with knowledge that’s two days old), performance remains stable. These results suggest that attacks crafted with slightly outdated knowledge of the model might have a lower effect on ML models. That indicates that continuously training the model might introduce an additional hurdle for attackers, as they would constantly need to obtain the updated gradients of the model, which can be a complex task, especially in the NIDS domain. Moreover, it is worth mentioning that this recovery of the model’s measures occurred after just one and two re-training sessions. Furthermore, these results were achieved without adversarial training, meaning that the re-training data didn’t contain adversarial samples [29]. For more detailed results, see Appendix B.
8.6 Interpretation of the Results
To visualize and gain an intuitive understanding of the data and results, we combine the visualization of Voronoi-based representations of multidimensional decision boundaries in scatterplots with the visualization of distances from data elements to the multidimensional boundary [55]. Data elements are reduced from 50 dimensions to 2 dimensions using the t-SNE (t-distributed Stochastic Neighbor Embedding) algorithm [56]. In other words, we show the evolution of the optimal multidimensional decision boundaries in 2D after every re-training session, as seen in Figure 16. Moreover, we also visualize the evolution of the data distribution after every re-training, as shown in Figure 17. As expected, the decision boundaries and training data distribution are always changing and evolving, as seen in Figures 16 and 17.
Figure 16: Decision Boundary Evolution using t-SNE
Because adversarial perturbations are highly aligned with a model’s weight vectors [59], when models undergo re-training, these weight vectors shift, rendering previously crafted perturbations misaligned with new decision boundaries. In real-world deployments where models update continuously, attackers face a timing problem: the model they analyze or query (maybe using side-channel querying, which takes time) during attack preparation may differ from the model they encounter at deployment time, a practical constraint rarely considered in research. However, the effectiveness of dynamic learning as a defense depends on multiple factors: model architecture, attack methodology, attacker knowledge, dataset characteristics, and the specific continual learning approach. While no flips, where adversarial samples remain misclassified after updates, or even negative flips may occur, where previously robust inputs become vulnerable after updates [60] (although Angioni et al. [60] tested negative flips using current attack knowledge, they could theoretically also occur with outdated adversarial knowledge) this complexity raises a fundamental question: if the system is dynamic and decision boundaries continuously shift, crafting adversarial perturbations becomes comparable to adding random noise. This challenges the practicality of employing extremely sophisticated attack methodologies when simpler attack vectors may prove more reliable and as effective in real-world scenarios.
9 Conclusion and Future Work
In this paper, we extend the research discussion on the practicality of adversarial attacks against NIDS by identifying the questionable practicality prerequisites required to carry out evasion adversarial attacks against NIDS using an attack tree and by presenting a taxonomy of practicality issues related to ML-based NIDS adversarial attacks. Our attack tree and taxonomy highlight several factors that could make numerous researched adversarial attacks impractical against real-world ML-based NIDSs. Additionally, we identified specific leaf nodes in our attack tree that demonstrate some practicality for real-world implementation and conducted a comprehensive review and exploration of these potentially viable attack approaches from the literature.
Figure 18: Adversarial Attacks Visualization
It is important to note that our work does not claim that adversarial attacks cannot harm ML-based NIDSs; on the contrary, we identified three potential routes where attackers can perform practical adversarial evasion attacks on NIDS. However, even these routes still have their own practicality limitations, which is why we find that the gap between research and real-world practicality is wide and deserves to be addressed. We also found that practical threat model assumptions are rare in current research papers, as most attack research still assumes generous threat model assumptions for the attacker. We further argue that this impractical attack research leads to overly expensive defense mechanisms that real-world ML-NIDSs might not actually need. Apruzzese et al. [3] argue that real-world attackers typically do not rely on gradients or conventional adversarial attack techniques; instead, they employ alternative simple yet effective approaches to evade ML models. If this is the case, then the current direction of ML security research needs a significant shift to address real-world challenges adequately. In the NIDS domain, we find that the most practical attacks are simple blind perturbations [42] that require absolutely no knowledge or querying and are very simple to generate. Nevertheless, there is no guarantee that these attacks will work consistently in practice.
We also explored the impact of adversarial attacks on dynamic ML-based NIDS, finding that continuous re-training, even without adversarial training, may limit the effect of some adversarial evasion attacks. It is important to note that simple blind perturbations will most likely not be affected by continuous re-training since they do not rely on gradients; however, these attacks remain inconsistent in their lethality and effectiveness. In future research, we plan to expand our attack taxonomy and apply it to network security applications beyond NIDS (e.g., spam detection). We will also explore how more sophisticated dynamic and continuous learning techniques will perform against adversarial attacks.
10 Acknowledgement
Appendix A Open Science
In alignment with open science policies, this research promotes transparency and reproducibility by:
  1. 1.
    Dataset: CSE-CIC-IDS2018 is a publicly available dataset.
  2. 2.
    Code Repository and Supplementary Materials: Detailed documentation, including instructions for reproducing experiments and analyzing results, as well as the code repository, will be made available upon request.
Appendix B Detailed Results
Pre-attack, as seen in Table 5 (a, b, and c), the NIDSs achieved high accuracy, precision, recall, and F1-score, indicating their effectiveness in detecting intrusions. However, after the FGSM attack, the performance dropped significantly for all three models. Among them, the SVM model experienced the highest drop, with its accuracy and F1 decreasing from 0.997 and 0.997 to 0.150 and 0, respectively. On the other hand, the ANN and CNN models had a decrease in accuracy and F1 of 15 to 30 percent. The reduction in accuracy and F1-measure was particularly prominent, indicating that the attacks successfully evaded the NIDSs’ detection. Nevertheless, continuous training improved the NIDSs’ resistance to FGSM attacks. On Day n+1 and Day n+2, we observed a significant recovery in accuracy, precision, recall, and F1-score.
Similarly, as shown in Table 5 (d to i), the results indicate that, like the FGSM attack, the performance of all three models significantly decreased after the PGD and BIM attacks. Among them, the SVM models experienced the most pronounced decline, while the ANN and CNN models exhibited a milder decrease in accuracy and F1. Continuous training also proved to enhance the NIDSs’ resistance to PGD and BIM attacks. On Day n+1 and Day n+2, we observed a substantial recovery in accuracy, precision, recall, and F1-score compared to Day n.
@@ Appendix B Detailed Results
Table 5: Effect of Continuous Training on FGSM, PGD and BIM Attacks against the ANN, SVM and CNN NIDSs, where Day n represents attacking before re-training, Day n+1 represents attacking one day after re-training, and Day n+2 represents attacking two days after re-training.